Circular No. 8348

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2004 K2 (McNAUGHT) A comet has been discovered by R. H. McNaught on CCD images ob-

tained with the 0.5-m Uppsala Schmidt telescope at Siding Spring (discovery observation below). Following posting on the NEO Confirmation Page, J. Young reported that CCD images obtained with the 0.6-m reflector at Table Mountain on May 25.5 UT show an 8′′–10′′ coma with no apparent central condensation and a straight, narrow tail 20′′–30′′ long in p.a. 252^{o}. CCD images taken by A. C. Gilmore and P. M. Kilmartin with the 0.6-m reflector at Mount John on May 25.7 show a circular 5′′ coma and no tail.

2004 UT  α_{2000}  δ_{2000}  Mag.
May 20.80223  23^{h}24^{m}13.^{s}21  −1^{o}31′06′′.4  17.7

The available astrometry, the following preliminary elliptical orbital elements, and an ephemeris appear on MPEC 2004-K61.

T = 2004 June 16.171 TT  ω = 180.^{o}362
e = 0.64243  Ω = 151.418 }2000.0
q = 1.59931 AU  i = 8.156
a = 4.47273 AU n^{o} = 0.104195 P = 9.46 years

COMETS C/2003 U6, U7, U8, U9 (SOHO) Further to IAUC 8341, K. Battams reports measurements for addi-

tional Kreutz sungrazing comets found on SOHO website C2 images (by X.-m. Zhou, C/2003 U7; X. Leprette, U8; M. Boschat, J. Sachs, and J. Danaher, U9); C/2003 U9 was also visible on C3 images. Sachs also found C/2003 U6 on SOHO website images. The reductions by R. Kisala and orbital elements by B. G. Marsden appear on MPEC 2004-K33.

Comet  2003 UT  α_{2000}  δ_{2000}
C/2003 U7 Oct. 23.064  13^{h}41.^{m}4  −12^{o}04′
C/2003 U8  25.788  13 56.1  −13 35
C/2003 U9  26.654  13 33.6  −15 21

SUPERNOVAE 2004bs, 2004bt, 2004bv Further to IAUC 8345, R. J. Foley et al. report the following unfiltered

CCD magnitude estimates from LOSS/KAIT images: SN 2004bs (IAUC 8341), Mar. 22.27 UT, [19.0; May 10.18, 16.7. SN 2004bt (cf. IAUC 8341), May 12.38, [18.5. SN 2004bv (IAUC 8344), May 16.50, [18.0.

2004 May 28  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green